OpenStudy (across):

t*y'-2*y=t^5*sin(2*t)-t^3+4*t^4 (PEMDAS notation) It's a differential equation. What's its solution? Let's see how good you guys are. :P

OpenStudy (across):

\[ty'-2y=t^5\sin(2t)-t^3+4t^4\]

OpenStudy (anonymous):

I assume you know how to do it?

OpenStudy (across):

I just want to corroborate my answer.

OpenStudy (anonymous):

Well Wolfram does it

OpenStudy (anonymous):

It seem little too advanced for me. I just started Fourier series

OpenStudy (across):

This is what I dribbled down to: \[y(t)=-(1/2)t^4\cos(2t)+(1/2)t^3\sin(2t)+1/4t^2\cos(2t)-t^3+2t^4+ct^2\]

OpenStudy (anonymous):

Sorry ...I am not that Good ..but I will bring you some Help

OpenStudy (anonymous):

Lemme see if I can work it out. It's first order linear ODE. Should be solvable using the integrating factor..

OpenStudy (across):

Yes, that's correct. I forgot to mention it's a linear differential equation. I used the integrating factor \[\mu(t)=t^{-2}\].

OpenStudy (anonymous):

Yep, that's what I have for the integrating factor also

OpenStudy (anonymous):

Yeah, I have what you got so far for y(t)

OpenStudy (across):

Thank goodness :) I just wanted to make sure I didn't mess up somewhere in between my calculations. Thank you! <3

OpenStudy (anonymous):

Nope, looks good. Messy, but I think it's right.